# of atoms in a sample of sulfur that weighs 12.02g (scientific notation)?
Answer #1

Molecular weigjt of sulpher( generallyexist as S8) = 256.5 g/mol

We know that 1 mol atom = 256.5 g S8 or molecular sulpher = 6.023 * 10e23 no of atoms.

So in 12.2 g of it = (12.2/256.5) * 6.023 * 10e23.

No of atom = 2.86 * 10e22

DO LIKE AND RATING..